OpenStudy (anonymous):

Solve for t. (t – 74) × 18 = 90 t = Solve for t. (t – 74) × 18 = 90 t = @Mathematics

OpenStudy (mathmagician):

t=90/18+74=79

OpenStudy (aroub):

18t-1332=90 18t= 1422 t=79

OpenStudy (anonymous):

thanks alot guys

OpenStudy (aroub):

no problem )

OpenStudy (anonymous):

t-74=5,t=79